# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ 6. Technical Lemmas
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be a subgraph of $\Gamma$ such that $\Gamma\setminus\Gamma^{\star}$ consists of $n$ $(1,d-1)$-fans;
  • $u\in U$ and $\mathcal{B}(u)$ be a ball in $\Gamma$ of $u$ of odd radius of at least $\max\{4h(d(n+1))+3,5\}$;
  • $\mathfrak{F}_{u}$ be an $(1,d-1)$-fan in $\Gamma^{\star}$ with the root $u$;
  • $\mathfrak{S}$ be an $(1,d-1)$-fan in $\Gamma\setminus\Gamma^{\star}$ with the root $u^{\perp}$;
  • $\mathfrak{M}$ be a perfect $(1,d)$-matching in the ball $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap\mathcal{B}(u)$.
If $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ with the witness
@@
but $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ does not satisfy Hall’s $d$-harem condition, then $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ satisfies Hall’s $d$-harem condition.
@@
We start the proof with several claims.
@@
Assume that $X\subset U^{\star}\setminus\{u\}$ does not satisfy Hall’s $d$-harem condition. If $X$ is not connected, then its neighbourhood is a disjoint union of the neighbourhoods of the connected components of $X$. Therefore without loss of generality we may assume that $X$ is connected. Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, there exists $u^{\prime}\in X$ such that distance between $u^{\prime}$ and $u$ is equal to $2$. On the other hand the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $((U^{\star}\setminus\mathfrak{F}_{u})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. Therefore $X\not\subseteq((U^{\star}\setminus\{u\})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. It follows that there exists $u^{\prime\prime}\in X\setminus\mathcal{B}(u)$. The distance between $u^{\prime}$ and $u^{\prime\prime}$ is at least $14h(d(n+1))+$ 1, i.e. $|X|\geq h(d(n+1))+1$. By the definition of $\hat{h}$,
@@
By this claim there is $X\subset V^{\star}\setminus\mathfrak{F}_{u}$ that does not satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$. Take the smallest connected one. Observe that $u$ must be in $N_{\Gamma^{\star}}(X)$. Indeed, otherwise $|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})}(X)|=|N_{\Gamma^{\star}}(X)|$. Since $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ we get a contradiction.
@@
For any vertex $v$ from $X$ the distance between $u$ and $v$ cannot exceed $12h(d(n+1))-$ 1.
@@
Since $\Gamma$ is bipartite, then distance between $u$ and any element of the set $X$ is odd. If for some $v\in X$ the distance between $u$ and $v$ is at least $12h(d(n+1))+$ 1, then $X$ has at least $h(d(n+1))$ elements. Since $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{\star}$, we have
@@
As a consequence of Claim 6.8, $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$.
@@
Since $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$ and $\mathfrak{M}$ satisfies the conditions of the perfect $(1,d)$-matching for all subsets of $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap(\mathcal{B}(u)\setminus\mathcal{S}(u))$, then
@@
If the distance between $u^{\perp}$ and $u$ is at least $22h(nd+1)+$ 2 then there exists an element of $X$ which is at distance at least $12h(nd+1)+$ 1 from $u$. Contradiction with Claim 6.8. ∎
We now see that $\{u,u^{\perp}\}\subset N_{\Gamma^{\star}\cup\mathfrak{S}}(X)$. Let us prove that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$. By Claim 6.7 all subsets of $U^{\star}\setminus\{u\}$ satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ and so in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ too. Therefore it remains to consider the case of subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ that contain $u^{\perp}$. Also note that the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. In particular trying to verify Hall’s $d$-harem condition in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for $Y\subset(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ we may assume that $Y\nsubseteq\mathcal{B}(u)$. As above we also assume that $Y$ is connected.
Let $u_{1}\in Y\setminus\mathcal{B}(u)$. The distance between $u$ and $u_{1}$ at least $44h(d(n+1))+$ 4. Since $u^{\perp}\in N_{\Gamma^{\star}\cup\mathfrak{S}}(X)\cap Y$, we see by Claim 6.9 that there is some $u_{2}\in Y$ at the distance $\leq 2h(d(n+1))$ from $u$. Then the distance between $u_{1}$ and $u_{2}$ is at least $4h(d(n+1))+4-(2h(d(n+1))+2)$, i.e at least $22h(d(n+1))+$ 2. Thus $|Y|>h(d(n+1))$, and
@@
It remains to show that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for any $Y\subset V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$. We may assume that $N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)\cap\{u,u^{\perp}\}\neq\emptyset$. Indeed, otherwise apply the assumption that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition. Also note again, that the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. In particular we may assume that $Y\nsubseteq(\mathcal{B}(u)\setminus\mathcal{S}(u))$. As above we only consider connected $Y\subset(V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$.
Take any element $v_{1}$ of $Y\setminus(\mathcal{B}(u)\setminus\mathcal{S}(u))$. It is clear that the distance from $u$ to $v_{1}$ is at least $34h(d(n+1))+$ 3. If $u\in N_{\Gamma^{\star}}(Y)$ then $Y$ contains a vertex at the distance $1$ from $u$. If $u\not\in N_{\Gamma^{\star}}(Y)$ then by Claim 6.9 $Y$ contains a vertex at the distance $12h(d(n+1))+$ 1 from $u$. In any case we can find $v_{2}\in Y$ at the distance at most $12h(d(n+1))+$ 1 from $u$. We see that the distance between $v_{1}$ and $v_{2}$ is at least $22h(d(n+1))+$ 2. Thus $Y$ has at least $h(d(n+1))$ elements and
@@
On the other hand using
@@
The last of technical lemmas concerns finding a witness of $c.e.H.h.c.(d)$ in a subgraph.
@@
  • $\mathfrak{M}$ be an $(1,d-1)$-matching in $\Gamma$ consisting of $\ell$ fans,
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be $\Gamma$ without $\mathfrak{M}$,
@@
Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition we may assume that $m>0$. First, we prove that $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for subsets of $U^{\star}$. The inequality $|X|\geq\widehat{h}(m)$ implies $|X|\geq h(m+\ell d)$, i.e.
@@
Therefore using $d\geq 2$ we see that when $|X|\geq\widehat{h}(m)$,
@@
Before stating the lemma proving claims 5.2, 5.6 we remind some of the relevant notation used at the $n+1$-st step of the construction.
@@
  • $\Gamma^{(n)\star}$ is a subgraph of $\Gamma^{(n)}$ obtained by removal of $(1,d-1)$-fans with roots belonging to the set $U^{(n)\perp}$;
  • $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$;
@@
If $U^{(n)\perp}=\emptyset$, then by Lemma 7.1 the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$. In particular, Claim 5.2 holds.
@@
Apply Lemma 6.10. In the first case view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$ in this lemma. The number $\ell$ from the lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by one, corresponding to the fan of $u_{n}$. Thus $\ell\leq 3n+1$. Since $\Gamma^{(n)\star}(-u_{n})$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $\Gamma^{\star}$. To compute the number corresponding to $\ell$ from the lemma we should take the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ together, then increase it by one, corresponding to the fan of $u_{n}$, and then subtract one, corresponding to the fan of $u_{j}^{\perp}$. Therefore $\ell\leq 3n$. Since $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
It is clear that both $\Gamma^{(n)\star}(-u_{n})$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ are highly computable bipartite graphs (assuming that $u^{\perp}_{j}$ is computed). Thus it is enough to show that one of these graphs satisfies Hall’s $d$-harem condition (then apply the claim). We know that the graph $\Gamma^{(n)\star}$ satisfies Hall’s $d$-harem condition. Let $\mathfrak{v}$ denote the only vertex from the set $\{v^{0}_{n,1},\ldots,v^{0}_{n,d}\}$ that belongs to $\Gamma^{(n)\star}(-u_{n})$. By Proposition 6.3 the choice of $u_{n},v^{0}_{n,1},\ldots,v^{0}_{n,d}$ ensures that $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ satisfies Hall’s $d$-harem condition as well. Let $\mathfrak{M}$ denotes the perfect $(1,d)$-matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$.
@@
Since the neighbourhood of any set is decomposed into a disjoint union of the neighbourhoods of connected subsets, applying the claim we see that if $\Gamma^{(n)\star}(-u_{n})$ does not satisfy Hall’s $d$-harem condition, this is witnessed by a finite connected $X\subset V^{(n)\star}(-u_{n})\cap(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$ which contains $\mathfrak{v}$. Using high computability of $\Gamma^{(n)}$ and finiteness of $\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$, we compute $X$ with these properties.
We will now prove that there exists $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. By the inequality $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$ it suffices to show that
@@
First, consider the case when $X=\{\mathfrak{v}\}$. The vertex $u_{\mathfrak{v}}$ either belongs to $U^{(n)\star}(-u_{n})$ or to $U^{(n)\perp}$. By inequality (1) and the fact that $U^{(n)\perp}$ consist of $(1,d-1)$-fans, in each case we have $|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|\geq d-1$. Moreover, the equality
@@
implies that $v_{u_{n}}\not\in\Gamma^{(n)}$. Indeed, if $v_{u_{n}}\in\Gamma^{(n)}$ then $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})$ by reflectedness. On the other hand equality (3) implies that $(u_{\mathfrak{v}},v_{u_{n}})\in\Gamma^{(n)\perp}$. Thus $(u_{n},\mathfrak{v})$ would be added to the matching $M$ at the first part of $n+1$-st step of the construction, i.e. $\mathfrak{v}\notin\Gamma^{(n)\star}(-u_{n})$, a contradiction. Indeed, if $v_{u_{n}}\in\Gamma^{(n)}$ then $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})$ (by reflectedness). On the other hand by the induction assumption, equality (3) implies that $(u_{\mathfrak{v}},v_{u_{n}})\in\Gamma^{(n)\perp}$. Thus $(u_{n},\mathfrak{v})$ would be added to the matching $M$ at the first part of $n+1$-st step of the construction, i.e. $\mathfrak{v}\notin\Gamma^{(n)\star}(-u_{n})$, a contradiction.
@@
This follows by $U^{(n)}$-reflectedness of $\Gamma^{(n)}$: $v_{u_{n}}$ is the only possible element adjacent to $u_{\mathfrak{v}}$ that does not have the left copy in $U^{(n)}(-u_{n})$. Additionally note that the equality
@@
$|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|>(d-1)-1\mbox{ and }|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|\geq\frac{1}{d}.$
@@
As we mentioned above this means that there is some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. Using high computability of $\Gamma^{(n)}(-u_{n})$ we can compute it.
Then $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|\geq 1$, and we have $|X|>d$ by the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$. Consider this case. Let $U_{X}:=\{u\in U:v_{u}\in X\}$ in $\Gamma^{(n)}(-u_{n})$. Applying the fact that $V^{(n)}$ is a subset of the copy of $U^{(n)}$ we arrive at two possibilities:

### v4

@@ 6. General Lemmas
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be a subgraph of $\Gamma$ such that $\Gamma\setminus\Gamma^{\star}$ consists of $n$ (1,d1)(1,d-1)-fans;
  • $u\in U$ and $\mathcal{B}(u)$ be a ball in $\Gamma$ of $u$ of odd radius $\geq\max\{4h(d(n+1))+3,5\}$;
  • $\mathfrak{F}_{u}$ be an (1,d1)(1,d-1)-fan in $\Gamma^{\star}$ with the root $u$;
  • $\mathfrak{S}$ be an (1,d1)(1,d-1)-fan in $\Gamma\setminus\Gamma^{\star}$ with the root $u^{\perp}$;
  • $\mathfrak{M}$ be an $(1,d)$-matching in the ball $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap\mathcal{B}(u)$ which satisfies the conditions of perfect $(1,d)$-matchings for all vertices that are in $\mathcal{B}(u)\setminus\mathcal{S}(u)$.
Assume that $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ with the witness
@@
If $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ does not satisfy Hall’s $d$-harem condition, then $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ already satisfies Hall’s $d$-harem condition.
@@
The proof contains several claims.
@@
Assume that $X\subset U^{\star}\setminus\{u\}$ does not satisfy Hall’s $d$-harem condition. If $X$ is not connected, then its neighborhood is a disjoint union of the neighborhoods of the connected components of $X$. Therefore, without loss of generality, we may assume that $X$ is connected. Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, there exists $u^{\prime}\in X$ such that the distance between $u^{\prime}$ and $u$ is equal to $2$. On the other hand, the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $((U^{\star}\setminus\mathfrak{F}_{u})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. Therefore $X\not\subseteq((U^{\star}\setminus\{u\})\cup\{u^{\perp}\})\cap\mathcal{B}(u)$. It follows that there exists $u^{\prime\prime}\in X\setminus\mathcal{B}(u)$. The distance between $u^{\prime}$ and $u^{\prime\prime}$ is at least $4h(d(n+1))+1$, i.e. $|X|\geq h(d(n+1))+1$. By the definition of $\hat{h}$,
@@
Assume that $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ does not satisfy Hall’s $d$-harem condition. By this claim, there is $X\subseteq V^{\star}\setminus\mathfrak{F}_{u}$ that does not satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$. We fix a connected one that is minimal. Observe that $u$ must be in $N_{\Gamma^{\star}}(X)$. Indeed, otherwise $|N_{(\Gamma^{\star}\setminus\mathfrak{F}_{u})}(X)|=|N_{\Gamma^{\star}}(X)|$. Since $\Gamma^{\star}$ satisfies $c.e.H.h.c.(d)$ we get a contradiction.
@@
For any vertex $v$ from $X$ the distance between $u$ and $v$ cannot exceed 2h(d(n+1))12h(d(n+1))-1.
@@
Since $\Gamma$ is bipartite, then distance between $u$ and any element of the set $X$ is odd. If for some $v\in X$ the distance between $u$ and $v$ is at least $2h(d(n+1))+1$, then $X$ has at least $h(d(n+1))$ elements. Since $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{\star}$, we have
@@
As a consequence of Claim 6.8, $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$. The following claim, together with the above observations, guarantees that $\{u,u^{\perp}\}\subset N_{\Gamma^{\star}\cup\mathfrak{S}}(X)$.
@@
Since $X\subset(\mathcal{B}(u)\setminus\mathcal{S}(u))$ and $\mathfrak{M}$ satisfies the conditions of perfect $(1,d)$-matchings for all subsets of $((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})\cap(\mathcal{B}(u)\setminus\mathcal{S}(u))$, then
@@
If the distance between $u^{\perp}$ and $u$ is at least $2h(nd+1)+2$ then there exists an element of $X$ which is at distance at least $2h(nd+1)+1$ from $u$. Contradiction with Claim 6.8. ∎
Let us prove that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$. By Claim 6.7 all subsets of $U^{\star}\setminus\{u\}$ satisfy Hall’s $d$-harem condition in $\Gamma^{\star}\setminus\mathfrak{F}_{u}$ and so in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ too. Therefore, it remains to consider the case of subsets of $(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ that contain $u^{\perp}$. Trying to verify Hall’s $d$-harem condition in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for $Y\subset(U^{\star}\setminus\{u\})\cup\{u^{\perp}\}$ we may assume that $Y\nsubseteq\mathcal{B}(u)$. Indeed, otherwise the matching $\mathfrak{M}$ already witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. As above, we also assume that $Y$ is connected.
Let $u_{1}\in Y\setminus\mathcal{B}(u)$. The distance between $u$ and $u_{1}$ is at least $4h(d(n+1))+4$. Since $u^{\perp}\in N_{\Gamma^{\star}\cup\mathfrak{S}}(X)\cap Y$, we see by Claim 6.9 that there is some $u_{2}\in Y$ at the distance $\leq 2h(d(n+1))$ from $u$. Then the distance between $u_{1}$ and $u_{2}$ is at least 4h(d(n+1))+4(2h(d(n+1))+2)4h(d(n+1))+4-(2h(d(n+1))+2), i.e at least $2h(d(n+1))+2$.
Figure 13.
Thus $|Y|>h(d(n+1))$, and
@@
It remains to show that Hall’s $d$-harem condition is satisfied in $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ for every $Y\subset V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$. We may assume that $N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)\cap\{u,u^{\perp}\}\neq\emptyset$. Indeed, otherwise apply the assumption that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition. Also note again that the matching $\mathfrak{M}$ witnesses Hall’s $d$-harem condition for all subsets of $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$ contained in $\mathcal{B}(u)\setminus\mathcal{S}(u)$. In particular, we may assume that $Y\nsubseteq(\mathcal{B}(u)\setminus\mathcal{S}(u))$. As above, we only consider connected $Y\subset(V\cap((\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S})$.
Take any element $v_{1}$ of $Y\setminus(\mathcal{B}(u)\setminus\mathcal{S}(u))$. It is clear that the distance from $u$ to $v_{1}$ is at least $4h(d(n+1))+3$. Recall our assumption that one of the elements $u,u^{\perp}$ is in $N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)$. If $u\in N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)$ then $Y$ contains a vertex at the distance $1$ from $u$. If $u^{\perp}\in N_{\Gamma^{\star}\cup\mathfrak{S}}(Y)$ then by Claim 6.9 $Y$ contains a vertex at the distance $2h(d(n+1))+1$ from $u$. In either case we can find $v_{2}\in Y$ at the distance at most $2h(d(n+1))+1$ from $u$. We see that the distance between $v_{1}$ and $v_{2}$ is at least $2h(d(n+1))+2$.
Figure 14.
Thus, $Y$ has at least $h(d(n+1))$ elements and
@@
On the other hand, using
@@
The final lemma of this section shows how to find a witness of $c.e.H.h.c.(d)$ in a subgraph.
@@
  • $\mathfrak{M}$ be an (1,d1)(1,d-1)-matching in $\Gamma$ consisting of $\ell$ fans,
  • $\Gamma^{\star}=(U^{\star},V^{\star})$ be the induced subgraph of $\Gamma$ without the vertices of $\mathfrak{M}$,
@@
Since $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, the case of $m=0$ is obvious. Assume that $m>0$. First, we prove that $\widehat{h}$ is a witness of $c.e.H.h.c.(d)$ for subsets of $U^{\star}$. The inequality $|X|\geq\widehat{h}(m)$ implies $|X|\geq h(m+\ell d)$, i.e.
@@
Therefore, using $d\geq 2$, we see that when $|X|\geq\widehat{h}(m)$,
@@
Before stating Lemma 7.1 (proving claims 5.2 and 5.6) we recall some notation used in case (2) of the $n+1$-st step of the construction.
@@
  • $\Gamma^{(n)\star}$ is a subgraph of $\Gamma^{(n)}$ obtained by removal of (1,d1)(1,d-1)-fans with roots belonging to the set $U^{(n)\perp}$;
  • $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$, and $u_{n}\in\Gamma^{(n)\star}$;
@@
If $U^{(n)\perp}=\emptyset$, then Lemma 7.1 implies that the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$. In particular, Claim 5.2 holds.
@@
Apply Lemma 6.10. In the first case, view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$ in this lemma. The number $\ell$ from the lemma is equal to the number of fans both in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and in $\Gamma^{(n)\perp}$, plus one, for the fan corresponding to $u_{n}$. At every step, at most one new fan appears in $\Gamma^{(n)\perp}$ and at most two new fans appear in $M_{n}\setminus\Gamma^{(n-1)\perp}$. Thus $\ell\leq 3n+1$. Since $\Gamma^{(n)\star}(-u_{n})$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case, view $\Gamma^{(0)}$ as $\Gamma$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $\Gamma^{\star}$. To compute the number corresponding to $\ell$ from the lemma, we should take the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ together, then increase it by one, corresponding to the fan of $u_{n}$, and then subtract one, corresponding to the fan of $u_{j}^{\perp}$. Therefore, $\ell\leq 3n$. Since $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
It is clear that both $\Gamma^{(n)\star}(-u_{n})$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ are highly computable bipartite graphs (assuming that $u^{\perp}_{j}$ is computed). Thus, it is enough to show that one of these graphs satisfies Hall’s $d$-harem condition (then apply the claim). We know that the graph $\Gamma^{(n)\star}$ satisfies Hall’s $d$-harem condition. Let $\mathfrak{v}$ denote the only vertex from the set $\{v^{0}_{n,1},\ldots,v^{0}_{n,d}\}$ that belongs to $\Gamma^{(n)\star}(-u_{n})$. By Proposition 6.3 the choice of $u_{n},v^{0}_{n,1},\ldots,v^{0}_{n,d}$ ensures that $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ satisfies Hall’s $d$-harem condition as well. For further arguments:
$\bullet$ let $\mathfrak{M}$ denote a perfect $(1,d)$-matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$.
@@
Since the neighborhood of any set is decomposed into a disjoint union of the neighborhoods of connected subsets, applying the claim, we see that if $\Gamma^{(n)\star}(-u_{n})$ does not satisfy Hall’s $d$-harem condition, then this is witnessed by a finite connected $X\subset V^{(n)\star}(-u_{n})\cap(\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n}))$ which contains $\mathfrak{v}$. Using high computability of $\Gamma^{(n)}$ and finiteness of $\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$, we compute $X$ with these properties; and, furthermore, it can be taken to be minimal. We will now prove the lemma into two steps.
$\bullet$ Step 1. There exists some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$.
This is the main part of the proof of the lemma. In fact we will prove that
@@
Then the existence of the required $u^{\perp}_{j}$ follows from the inequality $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$.
First, consider the case when $X=\{\mathfrak{v}\}$. The vertex $u_{\mathfrak{v}}$ either belongs to $U^{(n)\star}(-u_{n})$ or to $U^{(n)\perp}$. By inequality (1) and the fact that $U^{(n)\perp}$ consist of (1,d1)(1,d-1)-fans, in each case we have $|N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})|\geq d-1$. Moreover, if the equality
@@
holds, then $v_{u_{n}}\not\in\Gamma^{(n)}$. Indeed, if $v_{u_{n}}\in\Gamma^{(n)}$, then $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(u_{\mathfrak{v}})$ (by reflectedness). On the other hand, the existence of $\mathfrak{M}$ together with equality (3) implies that $(u_{\mathfrak{v}},v_{u_{n}})\in\Gamma^{(n)\perp}$. However, the first part of the $n+1$-st step of the construction implies that $(u_{n},\mathfrak{v})$ should be added to the matching $M_{n}$ already: the element $\mathfrak{v}$ plays the role of $v^{0}_{n,j}$ in case (2) of the first part. In particular, $\mathfrak{v}\notin\Gamma^{(n)\star}(-u_{n})$, a contradiction.
@@
This follows by $U^{(n)}$-reflectedness of $\Gamma^{(n)}$: $v_{u_{n}}$ is the only possible element adjacent to $u_{\mathfrak{v}}$ that does not have the left copy in $U^{(n)}(-u_{n})$. Additionally, note that the equality
@@
$|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|>(d-1)-1\mbox{ and, in particular, }|N_{\Gamma^{(n)}(-u_{n})}(\mathfrak{v})|\geq\frac{1}{d}.$
@@
As a result we have inequality (2) and the fact that there is some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$.
Consider this case when $X\not=\{\mathfrak{v}\}$. Then $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|\geq 1$ and by the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$, we have $|X|>d$.
Let $U_{X}:=\{u\in U:v_{u}\in X\}$ in $\Gamma^{(n)}(-u_{n})$. Applying the fact that $V^{(n)}$ is a subset of the right copy of $U^{(n)}$ we arrive at two possibilities: